OpenStudy (anonymous):

A triangular prism has a length of 12 inches. A larger similar triangular prism has a length of 18 inches. If 192 square inches of paper is used to wrap the smaller prism with minimum wastage, approximately how many square inches of paper will be required to wrap the larger prism?

OpenStudy (anonymous):

For two similar figures, the square of the ratio of two lengths is the ratio of two areas Lengths Areas a/b (a/b)^2 Since the ratio of 2 lengths is 18/12 = 3/2, the ratio of two area is (3/2)^2 = 9/4 9/4 = x/192 x = 9 x 192/4 = 432 square inches

OpenStudy (anonymous):

Thankyou so much

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Was 432 right?